# Efficient self-consistent learning of gate set Pauli noise

### v1

@@ I Introduction
Despite intensive recent research, some core aspects of the Pauli noise model are not yet well-understood. Consider learning the Pauli noise channel $\Lambda$ affecting a Clifford gate $\mathcal{G}$ (i.e., $\tilde{\mathcal{G}}=\Lambda\circ\mathcal{G}$) in the presence of unknown state-preparation-and-measurement (SPAM) noise. This is a standard task studied in multiple references [9, 11, 12, 13, 14, 15, 18, 19, 20]. When the ideal Clifford gate $\mathcal{G}$ equals to the identity, it has been shown that $\Lambda$ can be fully determined independent of the SPAM noise [9]; When $\mathcal{G}$ is a non-trivial Clifford gate, it is known that certain properties of $\Lambda$ (e.g., the average fidelity) can be learned independent of SPAM [10], but none of the protocols can completely and robustly learn $\Lambda$ in this case. In fact, it was pointed out [15, 25] that this is a fundamental limitation related to the model non-identifiability, also known as gauge ambiguity in the literature on gate set tomography [26]. Specifically, Chen, Liu et al[15] have developed a theory that completely characterize the degrees of freedom of a Pauli noise channel associated to any set of Clifford gates and provides a “gauge-consistent” algorithm to learn all the learnable degrees of freedom. The main limitation in Ref. [15] is that it considers the case that each $n$-qubit gate has a generic $n$-qubit Pauli noise channel, which has exponentially many parameters, and consequently the learning algorithm would take exponential time. In contrast, efficient parameterization of Pauli noise channels have been studied in the literature and applied in experiments [13, 14, 18, 16, 20, 27], but a good understanding about the gauge and learnable degrees of freedom in such models is yet to be obtained, which hinders the application of these protocols (specifically, learning such models self-consistently). It is thus natural to ask the following question: can one develop a Pauli noise learning protocol that is both self-consistent and efficient?
In this work, we address this question by introducing the framework of gate set Pauli noise learning: Given a set of multi-qubit Clifford gates and a noise model for each Pauli channel associated with the gates or SPAM, we obtain a complete characterization of the learnable degrees of freedom of that noise model, and provide a protocol to completely learn the learnable information whenever the true noise is realizable within the model. For a broad and experimentally relevant class of noise models with only polynomially many parameters, our learning protocol is efficient both for the number of experiments and for the classical computational complexity. Specifically, we show that all learnable noise parameters about the gates can be learned to relative precision for noise in this class in the realizable setting (i.e., all of the model assumptions hold exactly). The main technique we use comes from the algebraic graph theory. To demonstrate the power and flexibility of our framework, we apply it to the physically motivated models of spatially local or quasi-local noise. An illustration is given in Fig. 1. We give an analytic characterization of the learnable and gauge degrees of freedom for these noise models, and an explicit protocol for efficiently and self-consistently characterizing the noise.
@@ I.1 Summary of results
  1. (a)
    Learnability of gate set Pauli noise (Sec. III). We first set up a framework for gate set Pauli noise learning and provide general results on the model learnability. Focus on an $n$-qubit system. We assume any single qubit gates $\mathcal{U}$ can be implemented with negligible noise, while a fixed state preparation $\rho_{0}$, measurement $\{E_{j}\}_{j}$, and a set of multi-qubit Clifford gates denoted by $\mathfrak{G}$ can be implemented subject to Pauli noise channels. That is,
    $\tilde{\rho}_{0}=\Lambda^{S}(\rho_{0}),\quad\tilde{E}_{j}=\Lambda^{M}(E_{j}),\quad\tilde{\mathcal{G}}=\Lambda^{\mathcal{G}}\circ\mathcal{G},\quad\forall\mathcal{G}\in\mathfrak{G}.$ (1)
    We refer to the collection of all the above (noisy) operations as a (noisy) gate set, following the terminology of gate set tomography (GST) [26], and we refer to this noise model as the complete Pauli noise model, depicted in Fig. 1(b). The goal is to learn all the Pauli noise channels self-consistently, i.e., using only operations from the noisy gate set. It is known that fully determining the Pauli channels self-consistently is generally impossible [25, 15]. Indeed, consider the following gauge transformation on the noisy gate set,
    where $\mathcal{D}$ is some invertible linear map. It is not hard to see that any experimental outcome statistics remain unchanged under such transformation. Thus, two realizations of a noisy gate set related by a gauge transformation are indistinguishable. Meanwhile, one can carefully choose $\mathcal{D}$ (e.g., a depolarizing channel on a subset of qubits) such that the transformed noisy gate set will remain a valid Pauli noise model. That is, the single-qubit gates $\bigotimes_{i}\mathcal{U}_{i}$ remain noiseless, while $\Lambda^{S},\Lambda^{M},\Lambda^{\mathcal{G}}$ remain Pauli channels (possibly with different parameters). Now, consider a function of the gate set noise parameters. If the function value changes under certain gauge transformations, then the function is not self-consistently learnable. The first problem we aim to resolve is a characterization of the learnable and gauge degrees of freedom of the complete Pauli noise model.
    In this work, we encode all the model parameters into a linear space, called the parameter space. We then define learnable functions and gauge parameters, and show that they span two orthogonal complementary linear subspaces of the parameter space. Furthermore, by constructing a pattern transfer graph and using techniques from algebraic graph theory, we prove that the learnable space and gauge space are isomorphic to the cycle and cut space of the graph, respectively (Theorem 2). This gives a complete characterization of the learnability of the complete model. These results closely resemble Ref. [15] but are obtained from a different perspective.
    Since a generic Pauli noise channel contains exponentially many parameters, we often need some parameter-efficient noise ansatz such as some locality constraints of the Pauli channels, which contains a reduced number of noise parameters. We refer to such a noise model as a reduced Pauli noise model. Examples of reduced models with local and quasi-local constraints are shown in Fig. 1(c) and 1(d), respectively. To characterize the learnability of a reduced model, we introduce the notions of reduced parameter/learnable/gauge space similarly as for the complete model. Given certain mild assumptions of the noise ansatz, we characterize the reduced learnable and gauge space by making connections with the complete model, resolving the learnability of the reduced model (Theorem 3). These results answer the fundamental question of what is learnable from a Pauli noise model and what is not. They also serve as the basis for developing a self-consistent efficient learning algorithm.
  2. (b)
    Algorithms for learning gate set Pauli noise (Sec. IV). We introduce algorithms for learning a gate set Pauli noise model. We assume throughout that the noise is realizable, meaning that the true noise can be exactly expressed as a point in the parameter space. Given a reduced noise model, the goal is to learn everything within the reduced learnable space. We discuss two types of learning algorithms: The first one is tomography-like, consisting of experiments as simple as preparing certain initial product states, applying at most one noisy gate (or no gate at all), and measuring certain Pauli observables. These experiments can determine all learnable degrees of freedom in the large-sample limit; The second one aims at learning gate noise to relative precision, a desirable feature for many randomized benchmarking typed noise characterization protocols [28, 29, 9]. We prove that, for any learnable parameters of the noisy gates, one can always amplify the parameters by gate concatenation (Theorem 4), thus enabling a relative precision estimation of such parameters.
    Given that the reduced noise model is parameter-efficient, and that one can construct an appropriate “cycle basis” for the reduced learnable space, both of the above algorithms are efficient in the sense that they require at most polynomially many experiments to completely learn the reduced learnable space. However, efficiently constructing such a basis can require better knowledge about the specific noise ansatz. In the following two sections, we will apply the general results to some physically motivated noise models and obtain efficient learning algorithms.
  3. (c)
    Applications to the fully local noise model (Sec. V). We apply the general theory of learnability and learning algorithms to the fully local noise model. A fully local SPAM noise is a qubit-wise product Pauli channel, while a fully local gate noise is a Pauli channel acting non-trivially only on the gate support. Such a model has been considered in, e.g., ACES [13, 20, 27]. We prove that the reduced gauge space for a fully local noise is spanned by the single-qubit subsystem depolarizing gauges (SDG) (Theorem 5). We also construct a cycle basis for learning the fully local model using the first type of algorithm. To apply the second type of algorithm, we provide an efficient algorithm that can construct a cycle basis for the gate parameter space given that all noisy Clifford gates are $2$-qubit, which is the case for many existing quantum computing platforms (e.g., IBM Quantum [30]).
  4. (d)
    Applications to quasi-local noise model (Sec. VI). Going beyond the fully local model, we apply our theory to a quasi-local noise model that describes certain degrees of spatial correlation while remaining parameter-efficient. The quasi-local Pauli noise channels we consider are those whose Pauli eigenvalues factorize according to a Markov random field. We review useful properties of such noise model and discuss its connection and distinction with existing models studied in the literature [9, 14, 16]. We show that, under some validity assumptions of the quasi-local model, the reduced gauge space is spanned by the subsystem depolarizing gauges (SDG) consistent with the channel factorization (Theorem 6). Finally, a cycle basis for learning the quasi-local model using the first type of algorithm is also constructed.
  5. (e)
@@
    Case study: CZ gates with 1D topology (Sec. VII). To be even more concrete, we apply our theory to a concrete gate set where the Control-Z (CZ) gates are the only multi-qubit entangling gates. Such a gate set is adopted by state-of-the-art platforms [30]. We characterize the learnability of this noise model with more than one noise ansatz, and give efficient learning algorithms in each case.
I.2 Discussion and outlook
In what follows, we discuss the connection of our work with existing research, and point out a plethora of interesting open problems.

### v2

@@ I Introduction
Despite intensive recent study, some core aspects of Pauli noise learning have not been well-understood. In particular, a learning algorithm that is simultaneously self-consistent, comprehensive, and efficient is yet to be established. Here, self-consistency means using only operations that are noisy to learn about themselves, comprehensiveness means to learn all the information that is self-consistently learnable, and efficiency means the number of measurements and classical processing time scales favorable with the number of qubits. In this work, we build up a framework called gate set Pauli noise learning, give analytic characterization of the self-consistently learnable degrees of freedom, and develop an algorithm with the above three desirable features.
@@
Problem Setup. Consider an $n$-qubit quantum system. Suppose, by design, the following operations can be implemented on the platform: (1) Initialize the system to the ground state; (2) Apply a layer of $n$ single-qubit unitary gates; (3) Apply some multi-qubit Clifford gates gate from a finite set denoted by $\mathfrak{G}$; (4) Measure all qubits on the computational basis. We refer to the collection of all the above components as a gate set, following the language of gate set tomography (GST) [25], although we are also including noisy SPAM (state preparation and measurement). In practice, all the operations of a gate set can be noisy. Our goal is to characterize the noisy gate set self-consistently. That is, using only operations from the gate set to learn about themselves.
Instead of dealing with the most general form of noise, we focus on what is known as the Pauli noise model: We ignore the noise of any single-qubit gates, and assume the noise channels acting on the initial state, measurements, and any multi-qubit Clifford gate to be a (component-dependent) Pauli channel, which takes the following form,
$\Lambda(\rho)=\sum_{a\in{\sf P}^{n}}p_{a}P_{a}\rho P_{a}=\sum_{b\in{\sf P}^{n}}\lambda_{b}P_{b}\operatorname{Tr}(P_{b}\rho)/2^{n}.$ (1)
Here, $\mathsf{P}^{n}\mathrel{\mathop{:}}\nobreak\mkern-1.2mu=\{I,X,Y,Z\}^{\otimes n}$ is the $n$-qubit Pauli group (modulo phase). $\{p_{a}\}$ are the Pauli error rates while $\{\lambda_{b}\}$ are the Pauli fidelities or Pauli eigenvalues. We assume $\lambda_{b}>0$ throughout this work. Such a model can be ensured via randomized compiling under mild assumptions [23, 24]. The action of noise channel on the initial state $\rho_{0}$, POVM element $E_{j}$, and gate $\mathcal{G}\in\mathfrak{G}$ is given by
$\tilde{\rho}_{0}=\Lambda^{S}(\rho_{0}),\quad\tilde{E}_{j}=\Lambda^{M}(E_{j}),\quad\tilde{\mathcal{G}}=\Lambda^{\mathcal{G}}\circ\mathcal{G},$ (2)
respectively. A realization of the gate set Pauli noise model is specified by the involved Pauli channels $\{\Lambda^{S},\Lambda^{M},\Lambda^{\mathcal{G}}:\mathcal{G}\in\mathfrak{G}\}$. Since a generic Pauli channel contains exponentially many parameters, we often need to assume an efficient parametrization of the Pauli channels, which we call a noise ansatz. Examples of noise ansatzes include spatially local or quasi-local Pauli channels, as shown in Fig. 1(c) and (d), respectively. We refer to a Pauli noise model with a noise ansatz as a reduced model, and the one without a noise ansatz as a complete model. Throughout this work, we assume the true noise model is realizable. That is, the true noise model can be represented exactly as a point in the model space.
To characterize the noisy gate set means to learn all the Pauli noise channels with ansatz. However, it is generally impossible to fully determine those Pauli noise channels self-consistently [26, 16]. Indeed, consider the following gauge transformation [25] on the noisy operations,
where $\mathcal{D}$ is some invertible linear map. It is not hard to see that any experimental outcomes statistics remain unchanged under such transformation. Thus, two realizations of a noisy gate set related by a gauge transformation are indistinguishable. Meanwhile, one can carefully choose $\mathcal{D}$ (e.g., as a depolarizing channel on a subset of qubits) such that the transformed noisy gate set will remain a valid Pauli noise model within ansatz. That is, the single-qubit gates $\bigotimes_{i}\mathcal{U}_{i}$ remain noiseless, while $\Lambda^{S},\Lambda^{M},\Lambda^{\mathcal{G}}$ remain Pauli channels within the ansatz (possibly with different parameters).
Now, consider a function of the gate set noise parameters. If the function value changes under certain valid gauge transformations, then the function is not self-consistently learnable. This raises the first question of gate set Pauli noise learning as follows,
Problem 1.
How to classify the learnable and unlearnable functions of a gate set Pauli noise model?
This question is partially answered in [16] for the case that there is no constraint for the Pauli noise channels. When there are efficient ansatzes for the noise channels, the problem has remained open. After understanding what are the learnable degrees of freedom, the next natural problem is to find an efficient learning protocol, which can be summarized as,
Problem 2.
How to efficiently learn all the learnable information of a get set Pauli noise model?
In this work, we will establish a generic theoretic framework to address the above two problems. To demonstrate the power and flexibility of our framework, we will apply it to Pauli noise models with concrete gate set and noise ansatzes that are considered in the literature (e.g. [13, 15]) and resolve open problems therein. A detailed summary of our contributions is provided in the next section.
@@ I.1 Summary of Results
Learnability of gate set Pauli noise. Our approach to characterize the learnability (i.e. classification of learnable and unlearnable information) is to first encode all the noise parameters into a linear space. Let us start with the complete Pauli noise model (i.e. the one without a noise ansatz). Similar results have been obtained in [16] for this case, but here we will use a more formal way that can be extended to the reduced model later.
Define the logarithmic Pauli eigenvalues as $x_{b}\mathrel{\mathop{:}}\nobreak\mkern-1.2mu=-\log\lambda_{b}$. We collect all logarithmic Pauli eigenvalues (excluding the trivial ones corresponding to identity operator) from all Pauli noise channels into a ordered list denoted by $\bm{x}$, which can be viewed as a vector of a real linear space according to a standard basis. We denote the space $\bm{x}$ lives on as $X$, called the parameter space. An experiment is defined as a mapping $X\mapsto\mathbb{R}^{2^{n}}$ specified by a sequence of gates $\mathcal{C}$. It maps $\bm{x}$ to a measurement outcome probability distribution given by the Born rule, $\Pr(j)=\operatorname{Tr}[\tilde{E}_{j}\tilde{\mathcal{C}}(\tilde{\rho}_{0})]$, where the noisy realization of state, gates, and measurement is determined by $\bm{x}$. Now, we define a linear function $\bm{f}\in\mathcal{L}(X\mapsto\mathbb{R})\mkern-1.2mu=\mathrel{\mathop{:}}\nobreak X^{\prime}$ to be learnable if its value can be completely specified by a set of experiments [^1] One may wonder why we only consider linear functions of the logarithmic Pauli eigenvalues. This is because any experiments can be determined by some learnable linear functions, so they suffices to characterize the learnability. See Theorem 6 for more details.; We also define a vector $\mathfrak{d}\in X$ to be a gauge vector if the two lists of noise parameters $\bm{x}$ and $\bm{x}+\mathfrak{d}$ cannot be distinguished by any experiment, for all $\bm{x}\in X$. The linear subspace of all learnable linear function and all gauge vectors are denoted by $L$ and $T$, respectively. The problem then becomes characterizing $X$, $L$, and $T$.
For this purpose, we introduce a directed graph called the Pauli pattern transfer graph (PTG). This graph is first defined in [16] with some twist needed to fit into the current framework. The graph consists of $2^{n}$ nodes. One of which is a root node (also called a SPAM node), the others correspond to the $12^{n}-$ 1 non-zero $n$-bit strings, each denotes a Pauli pattern (the pattern of a $n$-qubit Pauli operator can be understood as its support); For each logarithmic Pauli eigenvalues, we assign a unique edge to the PTG: For gate noise parameters, each edge describe how a Clifford gate transform certain Pauli operators (where we only track the pattern); For each state preparation (or, measurement) noise parameter, we assign an edge from the root node to certain pattern node (or, from a pattern node to root). An example of PTG is given in Fig. 2.
Figure 2: The pattern transfer graph (PTG) for a two-qubit gate set $\mathfrak{G}=\{\mathrm{CZ}\}$. The gate set is depicted at the right of the graph. Here we define $\mathcal{G}=\mathrm{CZ}$ for notational simplicity. The black edges corresponds to gate noise parameters, while the gray edges corresponds to SPAM noise parameters. Multiple edges are shown as one edge with multiple labels.
PTG contains all information about a complete Pauli noise model. Intuitively, a path starting from and ending at the root node gives a combination of logarithmic Pauli eigenvalues that can be learned from an experiment. This intuition can be made rigorous using tools from algebraic graph theory [27, 28]. Informally, let the edge space $E$ to be the real linear space spanned by all edges of PTG. For each cycle (or, cut) on PTG, we can assign it with a vector of $E$ in a proper way (see Sec. III.2). The subspace spanned by all cycle (or, cut) vectors is defined as the cycle space $Z$ (or, the cut space $U$). It is known that $Z$ and $U$ are orthogonal complements to each other within $E$. Our main theorem is that, up to the natural identification between the parameter space $X$ and the edge space $E$, the learnable space $L$ is identical to the cycle space $Z$, while the gauge space $T$ is identical to the cut space $U$. This implies that $L$ and $T$ are also orthogonal complements to each other within $X$ [^2] Rigorously speaking, $L$ is a subspace of the dual space $X^{\prime}$. Here we identify $X$ with $X^{\prime}$ using the standard basis for defining $X$.
Next, we turn to study the learnability of a reduced model. A reduced model is specified by a tuple $(X_{R},\mathcal{Q})$. Here $X_{R}$ is the reduced parameter space, a linear space whose components are viewed as lists of the reduced noise parameters (denoted by $\bm{r}$). $\mathcal{Q}:X_{R}\mapsto X$ is the embedding map describing how the reduced parameters specifies the Pauli channels. Specifically, the logarithmic Pauli eigenvalues are given by $\bm{x}=\mathcal{Q}(\bm{r})$. We will assume $\mathcal{Q}$ is linear and injective. The notion of experiments on reduced model can be naturally induced from the definition for complete model. We can also similarly define the reduced learnable space $L_{R}$ (learnable functions of $X_{R}$) and the reduced gauge space $T_{R}$ (indistinguishable transformations within $X_{R}$). The problem is then to characterize $X_{R}$, $L_{R}$, $T_{R}$. The challenge here is that, it is not easy to construct a graph describing the reduced noise parameters as in the complete model. Our approach is to connect the reduced spaces back to the complete spaces, the use the characterization that we have already obtained. We proved that $L_{R}=\mathcal{Q}^{{\sf T}}(L)$$T_{R}=\mathcal{Q}^{-1}(T)$. In words, the reduced learnable space is the complete learnable space projected by $\mathcal{Q}^{\sf T}$, while the reduced gauge space is the preimage of the complete gauge space via $\mathcal{Q}$.
The relation between all the linear spaces introduced above can be summarized as follows.
Theorem 1 (Learnability, informal).
For any reduced model $(X_{R},\mathcal{Q})$ such that $\mathcal{Q}$ is linear and injective,
Reduced model: $\mathop{X_{R}}$ (reduced parameter) $=$ $\mathop{L_{R}}$ (reduced learnable) $\oplus^{\perp}$ $\mathop{T_{R}}$ (reduced gauge)
${}~{}\downarrow_{\mathcal{Q}}$ $\quad\uparrow_{\mathcal{Q}^{\sf T}}$ $\quad~{}~{}\uparrow_{\mathcal{Q}^{-1}}$
[…]
$\parallel$ $\parallel$ $\parallel$
[…]
The equality is up to the natural identification between $X$, $X^{\prime}$, and $E$. $\bigoplus^{\perp}$ means orthogonal complements.
Theorem 1 resolved Problem 1, giving a complete characterization about the learnable degrees of freedom for a Pauli noise model with general gate sets and noise ansatzes. They also serve as the basis for developing efficient self-consistent learning algorithms. Rigorous statements and detailed proofs for Theorem 1 are presented in Sec. III.
Efficient learning algorithms for gate set Pauli noise. We now consider Problem 2, how to efficiently and self-consistently learn a Pauli noise model. Thanks to Theorem 1, to learn a reduced Pauli noise model means to learn every functions of the reduced learnable space $L_{R}$. It suffices to find a basis for $L_{R}$ and learn the value of each basis function, as all the other functions can be decided by linearity. This can be done by first finding certain cycle basis for $E=L$ with each basis function learnable from some experiment, and then use the fact $L_{R}=\mathcal{Q}^{\sf T}(L)$ to pick up a subset of functions to form a basis for $L_{R}$. We give such a construction with the additional feature that each experiments uses at most one noisy gate, summarized as follows.
Theorem 2 (Simple experiment construction, informal).
There exists a set of $M=\dim(L_{R})$ experiments that can determine the value of any functions of $L_{R}$. Moreover, any experiment from this set either apply one gate from $\mathfrak{G}$ exactly once, or apply no gates from $\mathfrak{G}$ at all.
To show this, we first show that any cycle containing the root node exactly once (called a rooted cycle) can by learned by one experiment. Then notice that, thanks to the connectivity of PTG, there exists a rooted cycle basis where every cycle is of length 2 or 3, corresponding to experiment with no gate or exactly one gate from $\mathfrak{G}$. For an efficient noise ansatz, the number of reduced parameters should be at most polynomial in $n$, which means the number of experiments $M$ needed in Theorem 2 is also polynomial. The protocol can thus be efficiently executed. Details about this protocol are presented in Sec. IV.1.
In the literature of quantum noise learning, an important consideration is whether the noise parameter can be learned to relative precision. That is, given a sufficiently small infidelity parameter $r$ of a noisy gate $\tilde{\mathcal{G}}$, can we obtain an estimator $\hat{r}$ such that $|\hat{r}-r|\leq O(\varepsilon r)$ with high probability, by making a number of measurement that scales as $\varepsilon^{-2}\mathrm{polylog}(r^{-1})$? It is known that, if one can concatenate $m$ copies of $\tilde{\mathcal{G}}$ to amplify the infidelity to order $r^{m}$ for any $m$, this scaling can be achieved [29, Theorem 1]. We should that such noise amplification is indeed possible for gate set Pauli noise model. We have that,
Theorem 3 (Amplifying gate noise parameters, informal).
For any function that is a cycle vector and only supported on gate noise parameters, let $\lambda\mathrel{\mathop{:}}\nobreak\mkern-1.2mu=\exp(-\bm{f}(\bm{x}))$. Then, there exists a family of experiments $\{F_{m}\}_{m\in\mathbb{N}}$ and a Pauli observable, such that the expectation value of the $m$th experiment is given by $A\lambda^{m}$, where $A$ is an $m$-independent parameter.
Theorem 3 ensures one can achieve relative precision for a basis of learnable functions supported on the gate noise parameters, confirming that gate set Pauli noise learning has the same desirable feature as other quantum noise learning schemes such as randomized benchmarking [30]. Details about Theorem 3 are presented in Sec. IV.2.
Theorem 2 and 3 address Problem 2 from two different aspects, providing efficient experiment construction for gate set Pauli noise learning. The key is to find an appropriate basis for the reduced learnable space. One caveat is that, since the numbers of edges and nodes in PTG grow exponentially with the number of qubits $n$, an algorithm that first finds a cycle basis in PTG and then select a subset to construct a basis for $L_{R}$ will inevitably suffer from exponential run time. However, when the noise ansatzes have some structure, we can find such a basis analytically, thus circumventing the run time issue. In what follows, we demonstrate this point by applying our theory to concrete physically relevant noise ansatzes.
Applications to spatially-local noise ansatzes. We now apply the general theory of learnability and learning algorithms to concrete noise ansatzes. We will focus on spatially local or quasi-local noise model as an example. Such noise models have been widely studied in the literature of Pauli noise learning [9, 13, 20, 14, 15, 18, 31], but our results for the first time give a clear classification of the learnable and gauge degrees of freedom within such models.
We first consider the fully-local model, where the SPAM noise channels are assumed to be qubit-wise product channels, while the gate noise channels act non-trivially only within the gate’s support. This model is also sometimes known a crosstalk-free model and have been considered in, e.g.,  [13, 20, 31]. To understand the learnable and gauge spaces for this model, we introduce a family of gauge vectors called the subsystem depolarizing gauges (SDG), denoted by $\{\mathfrak{d}_{\nu}\}_{\nu\in 2^{[n]}}$, where $\mathfrak{d}_{\nu}$ represents a gauge transformation as in Eq. (3) with $\mathcal{D}$ chosen as a partially depolarizing channel acting on a subsystem of qubits $\nu$. One can show the SDGs forms a basis for the complete cut space $T$. Interestingly, we show that the subset of SDG’s acting on one qubits fully characterize the reduced gauge space $T_{R}$ for a fully-local model.
Theorem 4.
For any fully-local noise model $(X_{R},\mathcal{Q})$, its reduced gauge space satisfies,
$\mathcal{Q}(T_{R})=\mathrm{span}\{\mathfrak{d}_{s}:|s|=1\}.$ (4)
Since $\mathcal{Q}$ is injective, this allows one to uniquely determine $T_{R}$. Intuitively, any multi-qubit SDG transformation will violate the locality assumption of qubit-wise independent SPAM noise (and sometimes also local gate noise), and is thus an invalid gauge transformation. We also studied the more involved cases where either SPAM or gates has a fully-local noise assumption while the other does not. Even in that case, we can show that certain subset of SDG’s still spanned $\mathcal{Q}(T_{R})$. Detailed are presented in Sec. V.2.
Regarding learning protocols, the reduced learnable space $L_{R}$ and a set of $\dim(L_{R})$ experiments can be analytically written down. We further study the task of relative precision learning, which boils down to finding a cycle basis for the reduced gate noise parameters. Though we believe this is a hard problem in general, we give a computationally efficient procedure for the case where all Clifford gate from $\mathfrak{G}$ acts non-trivially on two qubits, a practical setting for most existing quantum computing architectures (e.g., [32]). These issues are addressed in Sec. V.3 and D.
Going beyond the fully local model, we apply our theory to a quasi-local noise model that describes certain degrees of spatial correlation while remains parameter-efficient. The quasi-local Pauli noise channels we consider are those whose Pauli eigenvalues factorize according to a factor graph [33]. We first introduce some notations that are needed for defining the quasi-local noise model, defined as follows
Definition 1.1.
for some $\{r_{b}\in\mathbb{R}:b\sim\Omega,b\neq I_{n}\}$ and a factor set $\Omega\subseteq 2^{[n]}$.
Here, $\Omega\subseteq 2^{[n]}$ is a factor set if it satisfies $\forall\nu\in\Omega$, $\forall\varnothing\subsetneq\mu\subseteq\nu$, $\mu\in\Omega$. We write $b\triangleleft a$ if $b_{i}\neq I\Rightarrow a_{i}=b_{i}$ for all $i\in[n]$. We write $b\sim\Omega$ if the support of $b$ belongs to $\Omega$. We show that, under certain “covariance” assumptions of the quasi-local model, the reduced gauge space is again characterized by a subset of SDGs that are consistent with the channel factorization,
Theorem 5.
(Informal) For any quasi-local noise model $(X_{R},\mathcal{Q})$, given that all Pauli channels are $\Omega$-local and is covariant for all gate in $\mathcal{G}$, then its reduced gauge space satisfies,
$\mathcal{Q}(T_{R})=\mathrm{span}\{\mathfrak{d}_{\nu}:\nu\in\Omega\}.$ (6)
Rigorous statements and detailed proofs are presented in Sec. VI, where we also review useful properties of such noise model and discuss its connection and distinction with existing models studied in the literature [9, 15, 14]. A cycle basis for the reduced learnable space $L_{R}$ is also analytically constructed for this case.
@@
Finally, in Sec. VII, we analyze a concrete gate set where the Control-Z (CZ) gates are the only multi-qubit entangling gates. We characterize the learnability of this noise model with more than one noise ansatzes (which are directly related to those that have been studied in the literature [15]) and give efficient learning algorithms in each case. These examples serve as an explicit recipe of gate set Pauli noise learning that is ready to be applied in current quantum devices.
I.2 Prior works
Learning the Pauli noise channel $\Lambda$ affecting a Clifford gate $\mathcal{G}$ (i.e., $\tilde{\mathcal{G}}=\Lambda\circ\mathcal{G}$) in the presence of unknown state-preparation-and-measurement (SPAM) noise is a standard task studied in multiple references [9, 11, 12, 13, 15, 16, 18, 19, 20]. When the ideal Clifford gate $\mathcal{G}$ equals to the identity, it has been shown that $\Lambda$ can be fully determined independent of the SPAM noise [9]; When $\mathcal{G}$ is a non-trivial Clifford gate, it is known that certain properties of $\Lambda$ (e.g., the average fidelity) can be learned independent of SPAM [10], but no existing protocols can completely and SPAM-robustly learn $\Lambda$ in this case. In fact, it was pointed out [16, 26] that this is a fundamental limitation related to the model non-identifiability, also known as gauge ambiguity in the literature on gate set tomography [25]. Specifically, [16] have developed a theory that completely characterize the degrees of freedom of a Pauli noise channel associated to any set of Clifford gates and provides a “gauge-consistent” algorithm to learn all the learnable degrees of freedom. The main limitation in [16] is that it considers the case that each $n$-qubit gate has a generic $n$-qubit Pauli noise channel, which has exponentially many parameters, and consequently the learning algorithm would take exponential time. In contrast, efficient parameterization of Pauli noise channels have been studied in the literature and applied in experiments [13, 15, 18, 14, 20, 31], but a good understanding about the gauge and learnable degrees of freedom in such models is yet to be obtained, which hinders the application of these protocols (specifically, how to learning such models self-consistently). Our work combines these two lines of research to achieve a self-consistent and efficient Pauli noise learning algorithm.
Our framework resembles ACES (i.e., averaged circuit eigenvalue sampling), a Pauli noise learning protocol proposed in [13] and further analyzed and implemented in [34] and [20]. Similar to our setting, ACES models every different Clifford gate as having a gate-dependent Pauli noise channel (with an efficient noise ansatz such as fully local noise), and aims at learning every Pauli eigenvalue of every Pauli noise channel. ACES constructs linear equations of the form $\bm{b}=A\bm{x}$, where $\bm{x}$ are the logarithmic Pauli eigenvalues, $\bm{b}$ are the logarithmic circuit eigenvalues which can be experimentally measured, and $A$ is the design matrix determined by a set of Clifford circuits and Pauli measurements. The original ACES effectively fixes a gauge by ignoring state preparation noise, and is thus able to construct a full column-rank $A$ using random Clifford gates. Within ACES’s framework, our work basically says that if SPAM noise parameters are included in $\bm{x}$, there is a maximum possible rank of $A$ equal to the number of learnable degrees of freedom, and we provide an explicit recipe for constructing such an $A$. Moreover, the explicit construction can learn gate parameters to relative precision; see Theorem 9. It is interesting to extend the statistical analysis for ACES [20] to our setting. It is also interesting to study optimal learning and experimental design within our framework [19, 35]. Conversely, it is interesting to study the ultimate sample complexity limit for learning gate set Pauli noise models. Such bounds for learning Pauli channels in a black-box setting have been previously investigated [17, 36, 37, 38, 39].
I.3 Future directions